OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 2, -8, 32, -128, ...

OpenStudy (anonymous):

Would the answer be a_n = 2 * (-4)^n or a_n = 2 * (-4)^n-1? I picked the first one but I just want to be sure.

OpenStudy (anonymous):

@iGreen

OpenStudy (saifoo.khan):

The second one.

OpenStudy (saifoo.khan):

Since that's what the formula states.

OpenStudy (michele_laino):

your sequence is a geometric sequence, with first term=2, and constant=-4, so: \[a _{n}=a _{1}q ^{n-1}=2*(-4)^{n-1}\]

OpenStudy (anonymous):

Thank you @Michele_Laino

OpenStudy (michele_laino):

Thank you! @sunshine201